find the equation of a line parallel to the graph of 2x-4y=5 that contains the point (0,3).
Answer by jim_thompson5910(35256)

2x-4y=C

2(0)-4(3)=C

0 - 12 = C

-12 = C

C = -12

So the equation of the parallel line in standard form is 2x-4y=-12

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------